OpenStudy (anonymous):

Subtracting a number from half its square gives a result of 13. Express the possible values of the number in simplest radical form.

OpenStudy (asnaseer):

\[\frac{x^2}{2}-x=13\]multiply both sides by 2 to get:\[x^2-2x=26\]therefore:\[x^2-2x-26=0\]solve this quadratic to get the two solutions for x.